# Discrete-state Continuous-time Diffusion for Graph Generation

### v1

@@ 3.4 Sampling Reverse Process
Given the parametric reverse rate matrix $\tilde{\mathbf{R}}_{\theta,t}(\mathcal{G},\bar{\mathcal{G}})$, the graph generation process can be implemented by two steps: (1) sampling the reference distribution $\pi_{\texttt{ref}}$ (i.e., $\pi_{f}$ for nodes and $\pi_{e}$ for edges) and (2) numerically simulating the CTMC from time $T$ to 00.
The exact simulation of a CTMC based on its rate matrix has been studied for a long time, e.g., [12, 13, 1]. However, their simulation strategies only allow one transition (e.g., one edge/node type change) per step, which is highly inefficient for graphs as the number of nodes and edges is typically large; once a(n) node/edge is updated, $\tilde{\mathbf{R}}_{\theta,t}$ requires recomputation. A practical approximation is to assume $\tilde{\mathbf{R}}_{\theta,t}$ is fixed, during a time interval $[t-\tau,t]$, i.e., delaying the happening of transitions in $[t-\tau,t]$ and triggering them all together at the time $t-\tau$; this strategy is also known as $\tau$-leaping [14, 7, 60]. DisCo adopts this sampling strategy.
We elaborate on $\tau$-leaping for transitions of node types; the transitions of edge types are similar. The rate matrix of the $i$-th node is fixed as $\tilde{\mathbf{R}}^{i}_{\theta,t}(f^{i},\bar{f}^{i})=\mathbf{R}_{t}^{i}(\bar{f}^{i},f^{i})\sum_{f_{0}^{i}}\frac{q_{t|0}(\bar{f}^{i}|f^{i}_{0})}{q_{t|0}(f^{i}|f^{i}_{0})}p^{\theta}_{0|t}(f^{i}|\mathcal{G}_{t})$, during $[t-\tau,t]$. According to the definition of rate matrix, in $[t-\tau,t]$, the number of transitions from $f^{i}$ to $\bar{f}^{i}$, namely $J_{f^{i},\bar{f}^{i}}$, follows the Poisson distribution, i.e., $J_{f^{i},\bar{f}^{i}}\sim\textrm{Poisson}(\tau\tilde{\mathbf{R}}^{i}_{\theta,t}(f^{i},\bar{f}^{i}))$. For categorical data (e.g., node type), multiple transitions in $[t-\tau,t]$ are invalid and meaningless. In other words, for the $i$-th node, if the total number of transitions $\sum_{\bar{f}^{i}}J_{f^{i},\bar{f}^{i}}>1$, $f^{i}$ keeps unchanged in $[t-\tau,t]$; otherwise, if $\sum_{\bar{f}^{i}}J_{f^{i},\bar{f}^{i}}=1$ and $J_{f^{i},s}>1$, i.e., there is exact $1$ transition, from $f^{i}$ to $s$, in $[t-\tau,t]$. The Appendix summarizes A step-by-step sampling algorithm in Algorithm 2.
@@
The sampling error of $\tau$-leaping is linear to $C_{\mathtt{err}}$ [7], where $C_{\mathtt{err}}$ is the approximation error of the reverse rate matrix: $\sum_{\mathcal{G}\neq\bar{\mathcal{G}}}\Big{|}\tilde{\mathbf{R}}_{t}(\mathcal{G},\bar{\mathcal{G}})-\tilde{\mathbf{R}}_{\theta,t}(\mathcal{G},\bar{\mathcal{G}})\Big{|}\leq C_{\mathtt{err}}$. Interested readers are referred to check Theorem 1 from [7]. Our Theorem 3.3 shows the connection between our training loss and $C_{\mathtt{err}}$, which further verifies the correctness of our training loss.
@@ 3.5 Model Instantiation
As mentioned in Section 3.3, the parametric backbone $p^{\theta}_{0|t}(\mathcal{G}_{0}|\mathcal{G}_{t})$ is a graph-to-graph mapping whose input is the noisy graph $\mathcal{G}_{t}$ and its output is the predicted denoised graph $\mathcal{G}_{0}$. There exists a broad range of neural network architectures. Notably, DiGress [62] uses a graph Transformer as the backbone model, a decent reference for our continuous-time framework. We name our model with the graph transformer (GT) backbone model as DisCo-GT and its detailed configuration is introduced in Appendix F.3. The main advantage of the GT is its long-range interaction thanks to the complete attention graph from the self-attention mechanism; however, the architecture is very complex and includes multi-head self-attention modules, leading to expensive computation.
@@
Beyond GTs, in this paper, we posit that a regular message-passing neural network (MPNN) [15] architecture should be a promising architecture for graph-to-graph mapping. It is recognized that the expressiveness of MPNNs might not be as good as GTs [28, 6], e.g., MPNN cannot model long-range interactions. However, in our setting, the absence of an edge is viewed as a special type of edge. In other words, the whole graph is complete; therefore, such a limitation of MPNN is naturally mitigated, which is verified by our empirical evaluations.
@@
Concretely, an MPNN-based graph-to-graph mapping is presented as follows, and our framework with MPNN is named DisCo-MPNN. Given a graph $\mathcal{G}=(\mathbf{E},\mathbf{F})$, where $\mathbf{E}\in\{1,\dots,a,a+1\}^{n\times n}$, $\mathbf{F}\in\{1,\dots,b\}^{n}$, we first transform both the matrix $\mathbf{E}$ and $\mathbf{F}$ into their one-hot representations $\mathbf{E}_{\texttt{OH}}\in\mathbb{R}^{n\times n\times(a+1)}$ and $\mathbf{F}_{\texttt{OH}}\in\mathbb{R}^{n\times b}$. Then, some auxiliary features (e.g., the # of connected components and the # of specific motifs) are extracted: $\mathbf{F}_{\texttt{aux}},\mathbf{y}_{\texttt{aux}}=\mathtt{Aux}(\mathbf{E}_{\texttt{OH}})$ to overcome the expressiveness limitation of MPNNs [10]. Here $\mathbf{F}_{\texttt{aux}}$ and $\mathbf{y}_{\texttt{aux}}$ are the node-level and global auxiliary features, respectively. Note that a similar auxiliary feature engineering is also applied in DiGress [62]. More details about the $\mathtt{Aux}$ can be found in Appendix E. Then, three multi-layer perceptrons (MLPs) are used to map node features $\mathbf{F}_{\texttt{OH}}\oplus\mathbf{F}_{\texttt{aux}}$, edge features $\mathbf{E}_{\texttt{OH}}$, and global features $\mathbf{y}_{\texttt{aux}}$ into a common hidden space as $\mathbf{F}_{\texttt{hidden}}=\mathtt{MLP}(\mathbf{F}_{\texttt{OH}}\oplus\mathbf{F}_{\texttt{aux}})$, $\mathbf{E}_{\texttt{hidden}}=\mathtt{MLP}(\mathbf{E}_{\texttt{OH}})$, $\mathbf{y}_{\texttt{hidden}}=\mathtt{MLP}(\mathbf{y}_{\texttt{aux}})$, where $\oplus$ is a concatenation operator. The following formulas present the update of node representations (e.g., $\mathbf{r}^{i}=\mathbf{F}(i,:)$), edge representations (e.g., $\mathbf{r}^{(i,j)}=\mathbf{E}(i,j,:)$), and global representation $\mathbf{y}$ in an MPNN layer, where we omit the subscript hidden if it does not cause ambiguity.
@@
where the edge embeddings are aggregated by mean pooling (i.e., $\frac{\sum_{j=1}^{n}\mathbf{r}^{(j,i)}}{n}$); the node pair embeddings are passed to edges by Hadamard product (i.e., $\mathbf{r}^{i}\odot\mathbf{r}^{j}$); edge/node embeddings are merged to the global embedding $\mathbf{y}$ via the PNA module [11]; Some FiLM modules [49] are used for the interaction between node/edge/global embeddings. More details about the PNA and FiLM are in Appendix E. In this paper, we name Eqs. (10), (11), and (12) on all nodes/edges together as an MPNN layer, $\mathbf{F},\mathbf{E},\mathbf{y}\leftarrow\mathtt{MPNN}(\mathbf{F},\mathbf{E},\mathbf{y})$. Stacking multiple MPNN layers can improve the capacity of the graph-to-graph backbone. In the final layer, two readout MLPs are used to project the node/edge embeddings into their input dimensions, $\mathtt{MLP}(\mathbf{F})\in\mathbb{R}^{n\times b}$ and $\mathtt{MLP}(\mathbf{E})\in\mathbb{R}^{n\times n\times(a+1)}$. These two sets of logits are output after wrapped with $\mathtt{softmax}$.
@@
Note that both the proposed MPNN layer and the GT layer from DiGress [62] use the PNA and FiLM to merge node/edge/global embeddings, but MPNN does not have a multi-head self-attention layer so that the computation overhead is lower.
@@ 3.6 Permutation Equivariance and Invariance
The proposed MPNN layer (Eqs. (10), (11), and (12)) is permutation-equivariant.
The auxiliary features from the $\mathtt{Aux}$ are also permutation-equivariant (see Appendix E). Thus, the whole MPNN-based graph-to-graph backbone $p^{\theta}_{0|t}$ is permutation-equivariant. Note that the GT-based backbone model from DiGress [62] is also permutation-equivariant whose proof is omitted as it is not our main contribution. Next, we show the permutation invariance of the forward and reverse rate matrices of DisCo.
@@
The forward rate matrix of DisCo is permutation-invariant if it is factorized as Eq. (5). The parametric reverse rate matrix of DisCo (Eq. (7)) is permutation-invariant whenever the graph-to-graph backbone $p^{\theta}_{0|t}$ is permutation-equivariant.
@@
The proposed training objective Eq. (9) is invariant to any permutation of the input graph $\mathcal{G}_{0}$ if the graph-to-graph backbone $p^{\theta}_{0|t}$ is permutation-equivariant.
@@ 4 Experiments
This section presents the effectiveness evaluation on the plain graph (Section 4.1) and molecule graphs (Section 4.2), an efficiency study (Section 4.3), and an ablation study (Section 4.4). Detailed experimental settings (Sections F.1-F.3), additional effectiveness evaluation (Section F.4), convergence study (Section F.5), and visualization of the generated graphs and the diffusion process (Section F.6) are in Appendix.
@@ 4.1 Plain Graph Generation
Datasets and metrics.
Datasets SBM, Planar [43], and Community [68] are used. The relative squared Maximum Mean Discrepancy (MMD) for degree distributions (Deg.), clustering coefficient distributions (Clus.), and orbit counts (Orb.) distributions (the number of occurrences of substructures with $4$ nodes), Uniqueness(%), Novelty(%), and Validity(%) are chosen as metrics. Results on Community [68] are reported in Table 8 in Appendix. Appendix F.2.2 provides more details about the datasets and metrics.
Baseline methods.
GraphRNN [68], GRAN [35], GG-GAN [30], MolGAN [8], SPECTRE [43], ConGress [62], DiGress [62] are chosen.
Results and analysis.
Table 1 shows
  • DisCo-GT can obtain competitive performance against the SOTA, DiGress, which is reasonable because both models share the same graph-to-graph backbone, graph Transformer. Note that DiGress’s performance in terms of Validity is not the statistics reported in the paper but from their latest model checkpoint  [^1] https://github.com/cvignac/DiGress/blob/main/README.md. In fact, we found it very hard for DiGress and DisCo-GT to learn to generate valid SBM/Planar graphs. These two datasets have only $200$ graphs, but sometimes only after $>10,000$ epochs training, the Validity percentage can be $>50\%$. Additionally, DisCo-GT provides extra flexibility during sampling by adjusting the $\tau$. This is important: our models can still trade-off between the sampling efficiency and quality even after the model is trained and fixed.
  • In general, DisCo-MPNN has competitive performance against DisCo-GT in terms of Deg., Clus., and Orb. However, its performance is worse compared to DisCo-GT in terms of Validity, which might be related to the difference in model expressiveness. Studying the graph-to-graph model expressiveness would be an interesting future direction, e.g., generating valid SBM and Planar graphs. As far as we know, no work exists in this direction.
Table 1: Generation performance on SBM and Planar datasets.
Model Deg.$\downarrow$ Clus.$\downarrow$ Orb.$\downarrow$ Unique $\uparrow$ Novel $\uparrow$ Valid $\uparrow$
SBM dataset
GraphRNN [68] 6.9 1.7 3.1 100.0 100.0 5.0
@@
4.2 Molecule Graph Generation
Dataset and metrics.
The datasets QM9 [52], MOSES [50], and GuacaMol [5] are chosen. For QM9, Uniqueness, Novelty, and Validity are reported in Table 2. For MOSES, metrics including Uniquess, Novelty, Validity, Filter, FCD, SNN, and Scaf are reported in Table 3. For GuacaMol, metrics including Valid, Unique, Novel, KL div, and FCD are reported in Table 4. More details about the datasets and the metrics are in Appendix F.2.3.
Baseline methods.
CharacterVAE [16], GrammarVAE [31], GraphVAE [56], GT-VAE [47], Set2GraphVAE [61], GG-GAN [30], MolGAN [8], SPECTRE [43], GraphNVP [42], GDSS [27], ConGress [62], DiGress [62], GRAPHARM [29],VAE [17], JT-VAE [24], GraphINVENT [45], LSTM [54], NAGVAE [33], and MCTS [23] are chosen.
Table 2: Generation performance (%) on QM9 dataset. V., U., and N. mean Valid, Unique, and Novel.
@@ 4.2 Molecule Graph Generation
DisCo-MPNN 98.9$\pm$0.7 98.7$\pm$0.5 68.7$\pm$0.2
Table 3: Generation performance on MOSES. VAE, JT-VAE, and GraphINVENT have hard-coded rules to ensure high validity.

### v2

@@ 3.4 Sampling Reverse Process
Given the parametric reverse rate matrix $\tilde{\mathbf{R}}_{\theta,t}(\mathcal{G},\bar{\mathcal{G}})$, the graph generation process can be implemented by two steps: (1) sampling the reference distribution $\pi_{\texttt{ref}}$ (i.e., $\pi_{f}$ for nodes and $\pi_{e}$ for edges) and (2) numerically simulating the CTMC from time $T$ to 00. The exact simulation of a CTMC has been studied for a long time, e.g., [16, 17, 1]. However, their simulation strategies only allow one transition (e.g., one edge/node type change) per step, which is highly inefficient for graphs as the number of nodes and edges is typically large; once a(n) node/edge is updated, $\tilde{\mathbf{R}}_{\theta,t}$ requires recomputation. A practical approximation is to assume $\tilde{\mathbf{R}}_{\theta,t}$ is fixed during a time interval $[t-\tau,t]$, i.e., delaying the happening of transitions in $[t-\tau,t]$ and triggering them all together at the time $t-\tau$; this strategy is also known as $\tau$-leaping [18, 8, 71], and DisCo adopts it.
We elaborate on $\tau$-leaping for transitions of node types; the transitions of edge types are similar. The rate matrix of the $i$-th node is fixed as $\tilde{\mathbf{R}}^{i}_{\theta,t}(f^{i},\bar{f}^{i})=\mathbf{R}_{t}^{i}(\bar{f}^{i},f^{i})\sum_{f_{0}^{i}}\frac{q_{t|0}(\bar{f}^{i}|f^{i}_{0})}{q_{t|0}(f^{i}|f^{i}_{0})}p^{\theta}_{0|t}(f^{i}|\mathcal{G}_{t})$, during $[t-\tau,t]$. According to the definition of rate matrix, in $[t-\tau,t]$, the number of transitions from $f^{i}$ to $\bar{f}^{i}$, namely $J_{f^{i},\bar{f}^{i}}$, follows the Poisson distribution, i.e., $J_{f^{i},\bar{f}^{i}}\sim\textrm{Poisson}(\tau\tilde{\mathbf{R}}^{i}_{\theta,t}(f^{i},\bar{f}^{i}))$. For categorical data (e.g., node type), multiple transitions in $[t-\tau,t]$ are invalid and meaningless. In other words, for the $i$-th node, if the total number of transitions $\sum_{\bar{f}^{i}}J_{f^{i},\bar{f}^{i}}>1$, $f^{i}$ keeps unchanged in $[t-\tau,t]$; otherwise, if $\sum_{\bar{f}^{i}}J_{f^{i},\bar{f}^{i}}=1$ and $J_{f^{i},s}=1$, i.e., there is exact $1$ transition, $f^{i}$ jumps to $s$. A step-by-step sampling algorithm (Algorithm 2) is in Appendix.
@@
The sampling error of $\tau$-leaping is linear to $C_{\mathtt{err}}$ [8], the approximation error of the reverse rates: $\sum_{\mathcal{G}\neq\bar{\mathcal{G}}}\Big{|}\tilde{\mathbf{R}}_{t}(\mathcal{G},\bar{\mathcal{G}})-\tilde{\mathbf{R}}_{\theta,t}(\mathcal{G},\bar{\mathcal{G}})\Big{|}\leq C_{\mathtt{err}}$. Interested readers are referred to Theorem 1 from [8]. Our Theorem 3.3 shows the connection between our training loss and $C_{\mathtt{err}}$, which further verifies the correctness of our training loss.
@@ 3.5 Model Instantiation
As mentioned in Section 3.3, the parametric backbone $p^{\theta}_{0|t}(\mathcal{G}_{0}|\mathcal{G}_{t})$ is a graph-to-graph mapping whose input is the noisy graph $\mathcal{G}_{t}$ and its output is the predicted denoised graph $\mathcal{G}_{0}$. There exists a broad range of neural network architectures. Notably, DiGress [73] uses a graph Transformer (GT) as $p^{\theta}_{0|t}$, a decent reference for our continuous-time framework. We name our model with the GT backbone as DisCo-GT and its detailed configuration is in Appendix F.3. The main advantage of the GT is its long-range interaction thanks to the complete self-attention graph; however, the architecture is very complex and includes multi-head self-attention modules, leading to expensive computation.
@@
Beyond GTs, in this paper, we posit that a regular message-passing neural network (MPNN) [19] should be a promising choice for $p^{\theta}_{0|t}(\mathcal{G}_{0}|\mathcal{G}_{t})$. It is recognized that the MPNNs’ expressiveness might not be as good as GTs’ [33, 7], e.g., in terms of long-range interactions. However, in our setting, the absence of an edge is viewed as a special type of edge and the whole graph is complete; therefore, such a limitation of MPNN is naturally mitigated, which is verified by our empirical evaluations.
@@
Concretely, an MPNN-based graph-to-graph mapping is presented as follows, and DisCo with MPNN backbone is named DisCo-MPNN. Given a graph $\mathcal{G}=(\mathbf{E},\mathbf{F})$, where $\mathbf{E}\in\{1,\dots,a,a+1\}^{n\times n}$, $\mathbf{F}\in\{1,\dots,b\}^{n}$, we first transform both the matrix $\mathbf{E}$ and $\mathbf{F}$ into one-hot embeddings $\mathbf{E}_{\texttt{OH}}\in\{0,1\}^{n\times n\times(a+1)}$ and $\mathbf{F}_{\texttt{OH}}\in\{0,1\}^{n\times b}$. Then, some auxiliary features (e.g., the # of specific motifs) are extracted: $\mathbf{F}_{\texttt{aux}},\mathbf{y}_{\texttt{aux}}=\mathtt{Aux}(\mathbf{E}_{\texttt{OH}})$ to overcome the expressiveness limitation of MPNNs [11]. Here $\mathbf{F}_{\texttt{aux}}$ and $\mathbf{y}_{\texttt{aux}}$ are the node and global auxiliary features, respectively. Note that a similar auxiliary feature engineering is also applied in DiGress [73]. More details about the $\mathtt{Aux}$ can be found in Appendix E. Then, three multi-layer perceptrons (MLPs) are used to map node features $\mathbf{F}_{\texttt{OH}}\oplus\mathbf{F}_{\texttt{aux}}$, edge features $\mathbf{E}_{\texttt{OH}}$, and global features $\mathbf{y}_{\texttt{aux}}$ into a common hidden space as $\mathbf{F}_{\texttt{hidden}}=\mathtt{MLP}(\mathbf{F}_{\texttt{OH}}\oplus\mathbf{F}_{\texttt{aux}})$, $\mathbf{E}_{\texttt{hidden}}=\mathtt{MLP}(\mathbf{E}_{\texttt{OH}})$, $\mathbf{y}_{\texttt{hidden}}=\mathtt{MLP}(\mathbf{y}_{\texttt{aux}})$, where $\oplus$ is a concatenation operator. The following formulas present the update of node embeddings (e.g., $\mathbf{r}^{i}=\mathbf{F}(i,:)$), edge embedding (e.g., $\mathbf{r}^{(i,j)}=\mathbf{E}(i,j,:)$), and global embedding $\mathbf{y}$ in an MPNN layer, where we omit the subscript hidden if it does not cause ambiguity:
$\mathbf{r}^{i}$ $\leftarrow\mathtt{FiLM}\bigg{(}\mathtt{FiLM}\bigg{(}\mathbf{r}^{i},\mathtt{MLP}\bigg{(}\sum_{j=1}^{n}\mathbf{r}^{(j,i)}/n\bigg{)}\bigg{)},\mathbf{y}\bigg{)},\ \ \mathbf{r}^{(i,j)}\leftarrow\mathtt{FiLM}\big{(}\mathtt{FiLM}(\mathbf{r}^{(i,j)},\mathbf{r}^{i}\odot\mathbf{r}^{j}),\mathbf{y}\big{)},$ (7)
$\mathbf{y}$ $\leftarrow\mathbf{y}+\mathtt{PNA}\big{(}\{\mathbf{r}^{i}\}_{i=1}^{n}\big{)}+\mathtt{PNA}\big{(}\{\mathbf{r}^{(i,j)}\}_{i,j=1}^{n}\big{)}.$ (8)
@@
The edge embeddings are aggregated by mean pooling (i.e., $\sum_{j=1}^{n}\mathbf{r}^{(j,i)}/n$); the node pair embeddings are passed to edges by Hadamard product (i.e., $\mathbf{r}^{i}\odot\mathbf{r}^{j}$); edge/node embeddings are merged to the global embedding $\mathbf{y}$ via the PNA module [12]; Some FiLM modules [57] are used for the interaction between node/edge/global embeddings. More details about the PNA and FiLM are in Appendix E. In this paper, we name Eqs. (7) and (8) on all nodes/edges together as an MPNN layer, $\mathbf{F},\mathbf{E},\mathbf{y}\leftarrow\mathtt{MPNN}(\mathbf{F},\mathbf{E},\mathbf{y})$. Stacking multiple MPNN layers leads to larger model capacity. Finally, two readout MLPs are used to project the node/edge embeddings into input dimensions, $\mathtt{MLP}(\mathbf{F})\in\mathbb{R}^{n\times b}$ and $\mathtt{MLP}(\mathbf{E})\in\mathbb{R}^{n\times n\times(a+1)}$, which are output after wrapped with $\mathtt{softmax}$.
@@
Both the proposed MPNN and the GT from DiGress [73] use the PNA and FiLM to merge embeddings, but MPNN does not have multi-head self-attention layers so that the computation overhead is lower.
@@ 3.6 Permutation Equivariance and Invariance
The proposed MPNN layer (Eqs. (7) and (8)) is permutation-equivariant.
The auxiliary features from the $\mathtt{Aux}$ are also permutation-equivariant (see Appendix E). Thus, the whole MPNN-based backbone $p^{\theta}_{0|t}$ is permutation-equivariant. Note that the GT-based backbone from DiGress [73] is also permutation-equivariant whose proof is omitted as it is not our contribution. Next, we show the permutation invariance of the rate matrices.
@@
The forward rate matrix of DisCo is permutation-invariant if it is factorized as Eq. (3). The parametric reverse rate matrix of DisCo ($\tilde{\mathbf{R}}_{\theta,t}$) is permutation-invariant whenever the graph-to-graph backbone $p^{\theta}_{0|t}$ is permutation-equivariant.
@@
The proposed training loss Eq. (6) is invariant to any permutation of the input graph $\mathcal{G}_{0}$ if $p^{\theta}_{0|t}$ is permutation-equivariant.
@@ 4 Experiments
This section includes: an effectiveness evaluation on plain graphs (Section 4.1) and molecule graphs (Section 4.2), an efficiency study (Section 4.3), and an ablation study (Section 4.4). Detailed settings (Sections F.1-F.3), additional effectiveness evaluation (Sections F.4, additional ablation study (Section F.5), convergence study (Section F.6), and visualization (Section F.7) are in Appendix. Our code is released  [^1] https://github.com/pricexu/DisCo.
@@ 4.1 Plain Graph Generation
Datasets and metrics. Datasets SBM, Planar [51], and Community [82] are used. The relative squared Maximum Mean Discrepancy (MMD) for degree distributions (Deg.), clustering coefficient distributions (Clus.), and orbit counts (Orb.) distributions (the number of occurrences of substructures with $4$ nodes), Uniqueness(%), Novelty(%), and Validity(%) are chosen as metrics. Details about the datasets, metrics, baselines (Section F.2.2), and results on Community (Table 8) are in Appendix.
Results. Table 1 shows the effectiveness evaluation on SBD and Planar from which we observe:
  • DisCo-GT can obtain competitive performance against the SOTA, DiGress, which is reasonable because both models share the graph Transformer backbone. Note that DiGress’s performance in terms of Validity is not the statistics reported in the paper but from their latest model checkpoint  [^2] https://github.com/cvignac/DiGress/blob/main/README.md. In fact, we found it very hard for DiGress and DisCo-GT to learn to generate valid SBM/Planar graphs. These two datasets have only $200$ graphs, but sometimes only after $>10,000$ epochs training, the Validity percentage can be $>50\%$. Additionally, DisCo-GT provides extra flexibility during sampling by adjusting the $\tau$. This is important: our models can still trade-off between the sampling efficiency and quality even after the model is trained and frozen.
  • In general, DisCo-MPNN has competitive performance against DisCo-GT in terms of Deg., Clus., and Orb. However, its performance is worse compared to DisCo-GT in terms of Validity, which might be related to the different model expressiveness. Studying the graph-to-graph model expressiveness would be an interesting future direction, e.g., generating valid Planar graphs.
Table 1: Performance (mean±$\pm$±std) on SBM and Planar datasets.
Dataset Model Deg.$\downarrow$ Clus.$\downarrow$ Orb.$\downarrow$ Unique $\uparrow$ Novel $\uparrow$ Valid $\uparrow$
SBM GraphRNN [82] 6.9 1.7 3.1 100.0 100.0 5.0
@@
Table 2: Performance (mean±$\pm$±std%) on QM9 dataset. V., U., and N. mean Valid, Unique, and Novel.
@@
4.2 Molecule Graph Generation
Dataset and metrics. The datasets QM9 [62], MOSES [58], and GuacaMol [6] are chosen. For MOSES, metrics including Uniquess, Novelty, Validity, Filter, FCD, SNN, and Scaf are reported in Table 3. For QM9, metrics include Uniqueness, Novelty, and Validity. For GuacaMol, metrics include Valid, Unique, Novel, KL div, and FCD. Details about the datasets, metrics, and baseline methods are in Appendix F.2.3.
Table 3: Performance on MOSES. VAE, JT-VAE, and GraphINVENT have hard-coded rules to ensure high validity.